OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). (sin x)(cos x) = 0

OpenStudy (ajprincess):

When a*b=0 a=0 or b=0. Similarly, here sinx=0 or cosx=0 Nw can find the solutions @needinghelpyo?

OpenStudy (anonymous):

no not really

OpenStudy (ajprincess):

\(\pi=180\)

OpenStudy (anonymous):

sinx cosx=1/2 sin2x

OpenStudy (anonymous):

0,90,180,270

OpenStudy (ajprincess):